calculate the number of grams of solute necessary to prepare the following solutions:


a. 25.0 m/ of a 0.200 N sulfuric acid solution in reactions that replace both hydrogen ions.


b. 50.0 m/ of a 1.25 N phosphoric acid solution in reactions that replace all three hydrogen ions.


1
Expert's answer
2022-05-16T17:02:04-0400

a)

1. using the following formula we determine the equivalent amount of solute:

nequivalent = CN * V = 0.2* 0.025= 0.005

2. determine the equivalent mass of sulfuric acid. To do this, we divide the molecular mass of the acid by its exchange Hydrogen ions number:

98/2= 49

3. To find the mass of an acid, we multiply its equivalent amount by its equivalent mass:

0.005*49=0.245 gr

Answer: 0.245 gr H2SO4

b)

1. using the following formula we determine the equivalent amount of solute:

nequivalent = CN * V = 1.25 * 0.05 = 0.0625

2. determine the equivalent mass of phosphoric acid. To do this, we divide the molecular mass of the acid by its exchange Hydrogen ions number:

98 / 3 = 32.6667

3. To find the mass of an acid, we multiply its equivalent amount by its equivalent mass:

0.0625 * 32.6667 = 2.04 gr

Answer: 2.04 gr H3PO4


ANSWER: a) 0.245gr H2SO4

b) 2.04gr H3PO4